# CasTex: Cascaded Text-to-Texture Synthesis via Explicit Texture Maps and Physically-Based Shading

### v1

@@ 6 Conclusion
In this work, we proposed an optimization-based texture synthesis approach based on cascaded diffusion models. We motivated the design with an analysis indicating the limitations of latent diffusion models. Our method improves upon recent optimization-based texture synthesis approaches in terms of automated metrics and user study. Even though our model benefits from using physically-based textures, we still observed baked-in lighting and occasional inconsistencies between a texture and the corresponding model geometry. Future work may address the baked-in lighting artifacts of the aforementioned limitations of Score Distillation Sampling with latent models.
@@ Acknowledgements
We would like to thank Roman Chulkevich for his assistance with the computing cluster and Ksenia Ushatskaya for her help with the visualization of the generated textures used in the paper teaser. Special thanks go to Maxim Kodryan, whose mere existence made this research journey considerably more enjoyable. This research was supported in part through computational resources of HPC facilities at HSE University.
@@ Appendix B Environment lights
Appendix C Texture Synthesis Results in Verying Setups
In this section, we provide textures obtained with different diffusion model combinations to illustrate the impact of model size and the role of super-resolution model.
@@ Appendix C Texture Synthesis Results in Verying Setups
We also provide additional illustrations to compare our method to the baselines.

### v2

@@ 6 Conclusion
In this work, we introduced CasTex, an optimization-based texture synthesis method that leverages cascaded pixel-space diffusion models with explicit texture parameterization. We motivated our design through an analysis highlighting the limitations of latent diffusion models. Our method demonstrates improved performance over recent optimization-based texture synthesis approaches, as measured by automated metrics and user studies. While CasTex benefits from physically-based textures, we still observe baked-in lighting and occasional inconsistencies between textures and the underlying geometry. Future work may address baked-in lighting artifacts or the aforementioned SDS limitations in latent-space models.
@@ Acknowledgements
The paper was prepared within the framework of the HSE University Basic Research Program and was supported in part through computational resources of HPC facilities at HSE University. This research was also supported by Yandex Education.
We would like to thank Roman Chulkevich and Viacheslav Kozyrev for their assistance with the computing cluster. We also thank Ksenia Ushatskaya for her help with the visualization of the generated textures used in the paper teaser. Special thanks go to Maxim Kodryan, whose mere existence made this research journey considerably more enjoyable.
@@ Appendix B Environment lights
Appendix C Texture Synthesis Results in Varying Setups
In this section, we present textures obtained with different diffusion model combinations to illustrate the impact of model size and the role of the super-resolution module. The results are shown in Figure 9. We also provide additional examples comparing our method with the baselines in Figure 10. Finally, Figure 11 shows a textured model after the first and the second stage of using our cascaded pixel diffusion, compared to the noisy output of latent diffusion.
@@ Appendix C Texture Synthesis Results in Varying Setups
Figure 11: Textured model after the first and the second stage of our cascaded pixel diffusion, compared to the noisy result from latent diffusion.
@@ Appendix E Visualization of PBR Texture Maps
Figure 13 further evaluates prompt‑conditioned control on a fixed mesh: we change only the material description and generate the corresponding PBR maps. The generator consistently shifts the distributions of metalness, roughness, and normal amplitude in a physically meaningful way, indicating that it learns material semantics rather than merely recoloring the surface. Metals remain the most sensitive case, where map statistics can still correlate with lighting; we attribute this to limited light map diversity and leave stronger lighting disentanglement for future work.
Figure 13: Prompt‑conditioned PBR textures demonstrating prompt‑level control over material properties. Top‑left (cork): high roughness variation, porous normal detail, near‑zero metalness. Top‑right (granite): stone‑like appearance from speckled albedo and fine normals; non‑metallic, mid–high roughness. Bottom‑left (soft‑touch plastic): uniform albedo, near‑zero metalness, satin roughness, low‑amplitude normals. Bottom‑right (steel): high metalness and lower roughness; residual noise and partially baked highlights remain—likely due to limited environment‑lighting variation; improving metal disentanglement is future work.